Two runners start running laps at the same time, from the same starting position. George runs a lap in 50s;Sally runs a lap in 30s. When will the runners next be side by side?
Answer by checkley77(12569)

HERE YOU NEED TO FIND THE LCD FOR THESE 2 TIMES
50=2*5*5
30=2*3*5
THE LCD=2*5*5*3=2*25*3=150 SECONDS THEY WILL MEET.